Determining the Optimal Product Mix with One Constrained Resource and a Sales Constraint

Casual Essentials, Inc. manufactures two types of team shirts, the Homerun and the Goalpost, with unit contribution margins of $5 and $15, respectively. Regardless of type, each team shirts must be fed through a stitching machine to affix the appropriate team logo. The firm leases seven machines that each provides 1,000 hours of machine time per year. Each Homerun shirt requires 6 minutes of machine time, and each Goalpost shirt requires 30 minutes of machine time.

Assume that a maximum of 50,280 units of each team shirts can be sold.

Required:

If required, round your answers to the nearest whole number.

1. What is the contribution margin per hour of machine time for each type of team shirts?

Contribution Margin
Homerun $50
Goalpost $30

2. What is the optimal mix of team shirts?

Optimal Mix
Homerun 50890??? units
Goalpost 3822??? units

3. What is the total contribution margin earned for the optimal mix?
$311,780???
Answer #1
  • Requirement 1 is CORRECT
  • Requirement 2
    >Total Hours available = 1000 hours per machine x 7 machines = 7000 hours

>Working

--Homerun has MORE contribution margin ($50) per hour than Goalpost ($30) per hour and hence HOMERUN should be given first preference to produce.

A

Total hours available

7000

B

Maximum units for Homerun

50280

C

Hours per unit required for Homerun

0.1

D = B x C

Total Hours required

5028

E = A- D

Remaining hours

1972

F

Hours per unit required for Goalpost

0.5

G = E/F

Total units for Goalpost

3944

>Requirement: Optimal Mix
Homerun = 50280 units
Goalpost = 3944 units

  • Requirement 3

Total Contribution margin = (50280 units x 0.1 hr x $ 50 per hour) + (3944 units x 0.5 hours x $ 30 per hour)
= 251400 + 59160
= $ 310,560 Answer